Circular No. 9024

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

2009 DD_{45}

R. P. Binzel, M. Birlan, and F. E. DeMeo, Paris Observatory, report on their 0.8- to 2.5-μm spectroscopic measurements of 2009 DD_{45} (cf. MPEC 2009-D80) on Mar. 2.6 UT using the NASA Infrared Telescope Facility 3-m reflector on Mauna Kea. Absorption bands revealed at 1 and 2 μm show the characteristics of the S-type class of minor planets. Using the average albedo value of 0.36 for small near-earth objects in this class (Delbo et al. 2003, Icarus 166, 116), and based on its H magnitude (25.4), the mean diameter is estimated to be 19 ± 4 m. The Apollo-type object passed only 0.000482 AU from the earth on Mar. 2.57 (cf. MPEC 2009-E10).

COMETS C/2008 V4, C/2008 V5, C/2008 V6 (SOHO) Further to IAUC 9022, additional near-sun presumed comets have been

found on SOHO website images, with their “discovery” observations tabulated below — all Kreutz sungrazers except for C/2008 V5 (non-group). C/2008 V4 was very faint (mag ∼ 8–8.5) and condensed. C/2008 V5 was perhaps very slightly diffuse but very small (mag ∼ 7.5–8). C/2008 V6 was small and condensed (mag ∼ 7.5).

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 V4 Nov. 7.063 14^{h}46.^{m}2 −17^{o}46′ C2 RM 2009-C46
C/2008 V5  8.660 15 04.1 −16 00 C2 JR 2009-C47
C/2008 V6  11.868 15 06.1 −18 57 C2 JR 2009-C47

COMET C/2007 N3 (LULIN)

Corrigendum. On IAUC 9023, line 12, for 267 ± 10 K, read 276 ± 10 K,

Further visual total-magnitude and coma-diameter estimates: Feb. 14.19 UT, 5.5, 10′ (B. H. Granslo, Fjellhamar, Norway, 7×50 binoculars); 20.52, 5.0, 20′ (M. Mattiazzo, Castlemaine, Victoria, Australia, 10×50 binoculars); 24.31, 4.7, 20′ (C. S. Morris, Fillmore, CA, U.S.A., 10×50 binoculars); 27.94, 4.9, 25′ (T. Karhula, Virsbo, Sweden, 8×40 binoculars); Mar. 3.14, 5.3, 25′ (J. J. Gonzalez, Leon, Spain, naked eye).

NOVA IN THE LARGE MAGELLANIC CLOUD 2009 G. Bianciardi, University of Siena, reports the following V magnitudes

for this nova (cf. IAUC 9019) from CCD images obtained remotely using a 0.35-m reflector near Sydney, N.S.W.: Feb. 25.717 UT, 14.50; 26.628, 14.72.

2009 March 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green